OpenStudy (anonymous):

For a circle of radius 8 feet, find the arc length s subtended by a central angle of 10 degrees.

jimthompson5910 (jim_thompson5910):

The angle of 10 degrees is 10/360 = 1/36 of a complete circle So the arc length will be 1/36 of the circumference

jimthompson5910 (jim_thompson5910):

Remember, the formula for the circumference is C = 2*pi*r

OpenStudy (goformit100):

^okay?

OpenStudy (anonymous):

That wasn't the answer I was looking for @jim_thompson5910 but thank you

jimthompson5910 (jim_thompson5910):

what did you get I cut myself short because I didn't want to do all of it

jimthompson5910 (jim_thompson5910):

and I wanted to let you do some and see what you get

OpenStudy (anonymous):

Well I'm given a choice between 4 answers and they are s=16/9 pi feet, s=8/9 pi feet, s=4/9 pi feet, and s=2/9 pi feet

jimthompson5910 (jim_thompson5910):

If the radius is 8 ft, then the circumference is what

jimthompson5910 (jim_thompson5910):

use C = 2*pi*r

jimthompson5910 (jim_thompson5910):

plug in r = 8 and leave the pi as it is and simplify

OpenStudy (anonymous):

50.27=C?

jimthompson5910 (jim_thompson5910):

leave the pi alone (ie don't replace it with 22/7 or 3.14 or anything like that) we leave the pi as it is because the answer has a pi in it

jimthompson5910 (jim_thompson5910):

C = 2*pi*r C = 2*pi*8 C = 2*8*pi C = 16*pi

jimthompson5910 (jim_thompson5910):

That's the exact distance around the circle the arc length s will be exactly 1/36th of this distance, so you just multiply 1/36 by 16pi and you get _____

OpenStudy (anonymous):

16/9 pi feet?

jimthompson5910 (jim_thompson5910):

16pi*(1/36) = 16pi/36 = ???

jimthompson5910 (jim_thompson5910):

basically you reduce 16/36 and then you tack on a "pi" at the end

OpenStudy (anonymous):

ohhhh my bad got ahead of myself there cx 4/9 pi feet

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

yay thank you n.n

jimthompson5910 (jim_thompson5910):

you're welcome